How many photons are produced in a laser pulse of 0.339 J at 505 nm? photons: 1.334 x10-29
Answer #1

E = hv

E = hc/λ

h is Planck's constant = 6.626⋅10−34.J s

c is the speed of light = 3 x 108.m s−1

λ =  wavelength

Let us convert the wavelength from nm to meter

wavelength = 505 nm=505 x 10−9  m.

Substitute the values and calculate the energy

E =   6.626⋅10−34 x 3 x 108 / 505 x 10−9 = 3.936 x 10−19  J

Let us calculate how many photons are needed.

0.339 J x 1 photons / 3.936 x 10−19  = 8.61 x 1017 photons

8.61 x 1017 photons are produced